A cell of emf 8.0 V has an internal resistance of 1.7 ȍ and is connected to a 4.9 ȍ resistor. Find the current in the resistor, the power dissipated in the resistor, and the terminal potential difference across the cell.


1
Expert's answer
2021-12-27T08:21:03-0500

emf:

"E=I(r+R)"

where r is internal resistance

"I=E\/(r+R)=8\/(1.7+4.9)=1.21" A


power dissipated in the resistor:

"P=I^2R=1.21^2\cdot4.9=7.20" W


terminal potential difference across the cell:

"U=Ir=1.21\cdot1.7=2.06" V